### v2

@@ 4.2 Interpreting NEOS
Figure 7 shows the contributions calculated using the keyword-based approach to determine $p_{t,i,j}$ for articles which contain the word “tariff”, respectively the corresponding German or French word. All the remaining articles which do not contain the keyword “tariff” are summarized in the category “Other”. In October 2024, the topic “Tariffs” began to contribute markedly negatively to NEOS. At that time, former President Donald Trump had pledged during his election campaign to impose additional tariffs if he were elected. The negative contribution peaked on President Trump’s “Liberation Day” in April 2025. It was also pronounced in August 2025 following the announcement of 39% additional tariffs on Swiss goods on 31 July.
Figure 7: Contribution to NEOS for topic “Tariffs” identified based on keywords.
@@
In Figure 9 we depict the contributions of the topics identified using the clustering approach. We show the clusters for which the absolute value of the monthly contribution exceeds 0.2 at least once and summarize the other clusters in “Other”. In line with the previous two contribution results, the word cloud in Figure 11 for Cluster 0 covers mainly topics related to the US’s tariffs. Cluster 1 covers technology related developments, in the economy in general or of particular firms and contributes mostly positive. An LLM-generated summary of the headlines closest to the cluster center per months is the following: Overall, these titles depict Switzerland’s dynamic role in the global tech ecosystem, emphasizing innovation, venture capital growth, international business expansion, and significant corporate moves (such as mergers and acquisitions). The Swiss tech sector seems to be benefiting from global trends, including the increasing prominence of start-ups, international investments, and the opportunities created by market instability elsewhere, especially in the US. The word cloud for Cluster 2 in Figure 11 shows mainly banking related topics. The most negative article from October 2025 highlights how fears of a banking crisis in the US also weighed on Swiss financial stocks. [^5] Our DUA with Swissdox@LiRI does not allow to publish the original news articles here as an illustration. Cluster 3 includes articles about various companies and their business situation, i.e. headlines like Quarterly results: Logitech exceeds expectations again in the second quarter (translated from German to English).
@@ 5 Sensitivity analysis
5.1 Evaluation of sentiment model
@@
To assess the quality of the predicted probabilities of the NEOS procedure (see Section 3.4), we conducted an experiment with two human labelers who were each assigned 75 news articles. These articles were randomly sampled based on their probability scores predicted in the NEOS procedure:
  • 50 articles from the 1st-5th quantile (lowest predicted probabilities), representing a negative economic outlook,
  • 50 articles from the 45th-55th quantile (around the median), representing a neutral economic outlook, and
  • 50 articles from the 95th-100th quantile (highest predicted probabilities), representing a positive economic outlook.
The labelers were instructed to classify the articles into three categories: negative, neutral, and positive economic outlook.
@@
The overall alignment between human and NEOS procedure labeling was 83%, with neutral articles showing the greatest divergence. Articles classified as neutral by the NEOS procedure were more frequently assigned to the positive or negative categories by human labelers. In contrast, the alignment for positive and negative articles was 93%. This is likely because the boundaries between articles classified as neutral by the NEOS procedure and those with a clear positive or negative outlook are more distinct. Since there is no strong bias in the disagreement for neutral articles, i.e., it is similar for positive and negative articles, this does not significantly affect NEOS.
5.2 Sensitivity to number of synthetic articles
@@
To assess whether 128 positive/128 negative synthetic articles are sufficient to ensure a stable sentiment score, we conducted an experiment where random subsamples of the synthetic articles were iteratively increased in size, and NEOS was computed for each subset. We found that using just 32 synthetic positive/negative articles resulted in high variance between the outcomes of different subsamples. However, as the number of synthetic articles increased, the variance decreased, with no further change after around 100 positive/negative synthetic articles. NEOS trained on subsets of around 100 synthetic positive/negative articles closely resembled the original NEOS results with 128 positive/128 negative synthetic articles. Increasing the number of synthetic articles further did not have any noticeable effect. Based on these findings, we conclude that a sample size of 128 positive/negative synthetic articles is sufficient for stable performance.
5.3 Leakage of future information in synthetic articles
@@
Training data up to April 2024 was used for Claude 3.5 Sonnet, i.e., the LLM which we use to generate the synthetic articles was trained on texts discussing past crises, like the EU debt crisis or the COVID-19 pandemic. It is not a priori clear whether our indicator would capture the sentiment of future crises similarly well compared to past crises which were known during training. To show that NEOS does not depend on such future information, we construct the indicator by training the sentiment model on the embeddings of 256 real news articles from 2012 instead of the synthetic articles. We choose the articles by selecting the ones with the highest, respectively lowest probability score from the originally trained logistic regression. The indicator obtained using real news articles from 2012 is shown together with the original monthly NEOS in Figure 12. The two indicators evolve in a similar way and have a correlation of 0.962 suggesting that our methodology does not rely on knowledge of future events in the synthetic articles. [^6] This type of leakage is not unique to our indicator. Compound survey indicators (like SECO-SEC, Swiss Economic Confidence) are based on multiple surveys. The composition of these indicators changes over time and their constituent subindices are selected based on historical correlations.
@@
Figure 12: The indicator obtained using news articles from 2012 and the one obtained using the synthetic articles highly correlate.
5.4 Relation of logistic regression to cosine similarity
@@
While this approach yields almost identical results, it is computationally much more costly because it requires computing $2\times N_{s}\times N$ similarity scores where $N_{s}$ is the number of positive/negative synthetic articles and $N$ the total number of real, relevant news articles.
@@
The synthetic positive and negative articles form two well-separated clusters in the embedding space (see Figure 4). Let $p=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}p_{i}$ and $n=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}n_{i}$ denote the average embeddings of the synthetic positive and negative examples. Then the cosine-similarity–based score
$\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\cos(x,p_{i})-\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\cos(x,n_{i})=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}x^{\top}(p_{i}-n_{i})=x^{\top}(p-n),$ (3)
shows that projecting $x$ onto the vector pnp-n measures whether $x$ aligns more with the positive or the negative cluster.
On the other hand, logistic regression predicts
$\Pr(y=1\mid x)=\sigma(x^{\top}w),\quad\sigma(z)=\frac{1}{1+e^{-z}},$ (4)
@@
$\sigma(x^{\top}w)\approx\text{const.}+x^{\top}w^{\prime}.$ (5)
Thus, up to scaling and shifting, logistic regression reduces to a linear score based on a direction $w^{\prime}$ that approximately corresponds to the separating vector pnp-n. In other words, the model uses a projection similar to a cosine-similarity-based approach. This provides additional affirmation that a simple linear method suffices in our setting.
@@ 6 Conclusion
Our indicator is transparent and interpretable, which is key for policymakers. It is based on a single data source and if required, each article’s influence on the indicator can be measured and the content can be retrieved for interpretation. We also presented approaches to allow for systematic interpretability over time and to allow detection of newly emerging topics.
Our proposed methodology offers a way of leveraging state-of-the-art LLMs even if data are proprietary, which is usually the case when working with full text news articles, and on-premise resources are limited. Moreover, the structure of our approach is modular and flexible. We can adjust the constituent embedding model or LLM to keep track of technical developments. The modularity of our approach also makes it straightforward to build new indicators to track developments in other variables or to use different data sources.
@@ A.1 Example of a prompt and output
As an example, to generate a synthetic news article covering financial markets we used the prompt below in Claude 3.5 Sonnet. All prompts were formulated in such a way that the generated articles convey a positive or negative outlook/assessment.
@@
Instead of using an LLM for text generation, we first experimented with short statements written by ourselves conveying positive or negative economic outlook, embedded them and used these embeddings to compute the indicator. We found, however, that our indicator improves significantly if the generated texts resemble real news articles in terms of length, style, structure etc. These findings are in line with gao:2022.
@@ A.2 Correlations with GDP growth
Table 2 benchmarks NEOS and its timelier versions against other survey-based indicators for economic outlook, and the lexicon-based approach introduced in Appendix A.3. The table shows the average correlations of the indicators with real year-on-year (yoy) and quarter-on-quarter (qoq) GDP growth for different lags over the period Q1 1999 until Q1 2025. Since GDP is a quarterly time series, we transform all monthly indicators to quarterly time series by taking three-month averages over the respective quarter. For instance, a lag of three quarters means that we compare GDP growth at time $t$ with the respective indicator at time t3t-3.
NEOS correlates with GDP growth in a similar way to the other indicators and outperforms for certain lags. The correlations with quarter-on-quarter GDP growth are much lower for all indicators. [^7] If qoq correlations with GDP growth are positive over multiple lags, yoy correlations will be higher by construction. Yoy GDP growth contains not only contemporaneous information but also information up to 12 months ago. NEOS shows the highest correlation for a lag of one quarter with respect to year-on-year GDP. The contemporaneous correlations (lag 0) can still be interpreted as a lead of the indicator since GDP is released with a delay of two months. [^8] The flash estimate is delayed by less than two months ($t+45$ days) Note that Table 2 shows ex post comparisons and does not account for actual data availability of the other indicators in the past and GDP revisions. Not all indicators are released immediately at the end of the month and historically, some indicators are only available quarterly.
Table 2: Analysis from Q1 1999 until Q1 2025
@@
NEOS 0.62 0.68 0.64 0.50 0.20 0.65 0.53
NEOS (first 7 days in month) 0.59 0.62 0.60 0.50 0.19 0.61 0.50
NEOS (first 14 days in month) 0.60 0.66 0.63 0.51 0.19 0.63 0.52
NEOS (first 21 days in month) 0.59 0.67 0.64 0.52 0.21 0.63 0.53
Lexicon-based approach 0.56 0.56 0.46 0.35 0.11 0.56 0.41
Manufacturing PMI for Switzerland 0.75 0.73 0.60 0.39 0.11 0.74 0.51
[…]
SECO Consumer Sentiment Index 0.51 0.45 0.38 0.24 0.00 0.48 0.32
@@
NEOS 0.45 0.23 0.08 0.11 -0.07 0.34 0.16
NEOS (first 7 days in month) 0.48 0.16 0.13 0.11 -0.07 0.32 0.16
NEOS (first 14 days in month) 0.48 0.20 0.12 0.10 -0.08 0.34 0.16
NEOS (first 21 days in month) 0.46 0.22 0.10 0.12 -0.09 0.34 0.17
Lexicon-based approach 0.35 0.16 -0.01 0.09 -0.04 0.26 0.11
Manufacturing PMI for Switzerland 0.41 0.16 0.08 0.02 -0.09 0.28 0.12
[…]
SECO Consumer Sentiment Index 0.32 0.01 0.07 0.00 -0.09 0.16 0.06
RMSE ratios with Diebold–Mariano p-values in parentheses. Best RMSE values in bold.
@@
a Correlations computed from Q1 2014 until Q1 2025 due to availability of the indicator.
b Correlations computed from Q2 2009 until Q1 2025 due to availability of the indicator.
@@
A.3 Lexicon-based approach to measure sentiment from news articles

### v3

@@ 4.5 Interpreting NEOS
Figure 8 shows the contributions calculated using the keyword-based approach to determine $p_{t,i,j}$ for articles that contain the German or French word for “tariff”. All the remaining articles that do not contain the keyword “tariff” are summarized in the category “other”. In October 2024, the topic “tariffs” began to contribute markedly negatively to NEOS. At that time, former President Donald Trump had pledged during his election campaign to impose additional tariffs if he were elected. The negative contribution peaked on President Trump’s “Liberation Day” in April 2025. It was also pronounced in August 2025 following the announcement of 39% additional tariffs on Swiss goods on July 31.
Figure 8: Contribution of the topic “tariffs” to NEOS as identified based on keywords.
@@
In Figure 10 we depict the contributions of the topics identified using the clustering approach. We show the clusters for which the absolute value of the monthly contribution exceeds 0.2 at least once and summarize the other clusters in “other”. In line with the previous two contribution results, the word cloud in Figure 12 for Cluster 0 covers mainly topics related to the US’s tariffs. Cluster 1 covers technology related developments, in the economy in general or of particular firms and contributes mostly positive. An LLM-generated summary of the headlines closest to the cluster center per month is the following: Overall, these titles depict Switzerland’s dynamic role in the global tech ecosystem, emphasizing innovation, venture capital growth, international business expansion, and significant corporate moves (such as mergers and acquisitions). The Swiss tech sector seems to be benefiting from global trends, including the increasing prominence of start-ups, international investments, and the opportunities created by market instability elsewhere, especially in the US. The word cloud for Cluster 2 in Figure 12 shows mainly banking-related topics. The most negative article from October 2025 highlights how fears of a banking crisis in the US also weighed on Swiss financial stocks. [^5] Our DUA with Swissdox@LiRI does not allow the original news articles to be published here as an illustration. Cluster 3 includes articles about various companies and their business situation, i.e., headlines such as Quarterly results: Logitech exceeds expectations again in the second quarter (translated from German to English).
@@ 5 Sensitivity analysis
5.1 Evaluation of the sentiment model
@@
To assess the quality of the predicted probabilities of the NEOS procedure (see Section 3.4), we conducted an experiment with two human labelers who were assigned the same 150 news articles each. We randomly sample 50 articles each from three parts of the overall indicator distribution: 50 articles from the 1st-5th quantile (the lowest scores representing a negative outlook), 50 articles from the 45th-55th quantile (scores near the median representing a neutral outlook), and 50 articles from the 95th-100th quantile (the highest scores representing a positive outlook). The labelers were instructed to classify the articles into three categories: negative, neutral, and positive economic outlook.
@@
The average alignment between the human and NEOS procedure labeling was 81%, while the alignment between the two human labelers was 79%. This implies that the labels generated using the NEOS procedure are at least as reliable as those produced by human labelers, and potentially more consistent. Neutral articles show the greatest divergence. Articles classified as neutral by our method were more frequently assigned to the positive or negative categories by human labelers. In contrast, the alignment for positive and negative articles was 89%.
5.2 Sensitivity to the number of synthetic articles
@@
To assess whether 128 positive/128 negative synthetic articles are sufficient to ensure a stable sentiment score, we conduct an experiment where random subsamples of the synthetic articles are iteratively increased in size, and NEOS is computed for each subset. We found that using just 32 synthetic positive/negative articles resulted in high variance between the outcomes of different subsamples. However, as the number of synthetic articles increased, the variance decreased, with no further change after approximately 100 positive/negative synthetic articles. NEOS trained on subsets of approximately 100 synthetic positive/negative articles closely resembled the original NEOS results with 128 positive/128 negative synthetic articles. Increasing the number of synthetic articles further did not have any noticeable effect. Based on these findings, we conclude that a sample size of 128 positive/negative synthetic articles is sufficient for stable performance.
5.3 Leakage of future information in the synthetic articles
@@
Training data up to April 2024 was used for Claude 3.5 Sonnet, i.e., the LLM we use to generate the synthetic articles was trained on texts discussing past crises, such as the EU debt crisis or the COVID-19 pandemic. It is not a priori clear whether our indicator would capture the sentiment of future crises similarly well compared to past crises that were known during training. To show that NEOS does not depend on such future information, we construct the indicator by training the sentiment model on the embeddings of 256 real news articles from 2012 instead of the synthetic articles. We choose the articles by selecting the ones with the highest and lowest probability score from the originally trained logistic regression. The indicator obtained using real news articles from 2012 is shown together with the original monthly NEOS in Figure 13. The two indicators evolve in a similar way and have a correlation of 0.962 suggesting that our methodology does not rely on knowledge of future events in the synthetic articles. [^6] This type of leakage is not unique to our indicator. Compound survey indicators (such as SECO-SEC and Swiss Economic Confidence) are based on multiple surveys. The composition of these indicators changes over time and their constituent subindices are selected based on historical correlations.
@@
Figure 13: High correlation between the indicator obtained using news articles from 2012 and the one obtained using the synthetic articles.
5.4 Relation of the logistic regression to cosine similarity
@@
While this approach yields almost identical results, it is computationally much more costly because it requires computing $2\times N_{s}\times N$ similarity scores, where $N_{s}$ is the number of positive/negative synthetic articles and $N$ the total number of real, relevant news articles.
@@
The synthetic positive and negative articles form two well-separated clusters in the embedding space (see Figure 4). Let $p=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}p_{i}$ and $n=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}n_{i}$ denote the average embeddings of the synthetic positive and negative examples, respectively. The cosine-similarity–based score
$\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\cos(x,p_{i})-\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}\cos(x,n_{i})=\frac{1}{N_{s}}\sum_{i=1}^{N_{s}}x^{\top}(p_{i}-n_{i})=x^{\top}(p-n),$
shows that projecting $x$ onto the vector pnp-n measures whether $x$ aligns more with the positive or negative cluster.
On the other hand, the logistic regression predicts
$\Pr(y=1\mid x)=\sigma(x^{\top}w),\quad\sigma(z)=\frac{1}{1+e^{-z}},$
@@
$\sigma(x^{\top}w)\approx\text{const.}+x^{\top}w^{\prime}.$
Thus, up to scaling and shifting, the logistic regression reduces to a linear score based on a direction $w^{\prime}$ that approximately corresponds to the separating vector pnp-n. In other words, the model uses a projection similar to a cosine-similarity-based approach. This provides additional affirmation that a simple linear method suffices in our setting.
@@ 6 Conclusion
Our indicator is transparent and interpretable, which is key for policymakers. It is based on a single data source and if required, each article’s influence on the indicator can be measured and the content can be retrieved for interpretation. We also presented approaches to allow for systematic interpretability over time and to allow the detection of newly emerging topics.
Our proposed methodology offers a way of leveraging state-of-the-art LLMs even if data are proprietary, which is usually the case when working with full text news articles, and on-premises resources are limited. Moreover, the structure of our approach is modular and flexible. We can adjust the constituent embedding model or LLM to keep track of technical developments. The modularity of our approach also makes it straightforward to build new indicators to track developments in other variables or to use different data sources.
@@ A.1 Example of a prompt and output
As an example, to generate a synthetic news article covering financial markets, we used the prompt below in Claude 3.5 Sonnet. All prompts were formulated in such a way that the generated articles convey a positive or negative outlook/assessment.
@@
Instead of using an LLM for text generation, we first experimented with short statements written by ourselves conveying positive or negative economic outlook, embedded them and used these embeddings to compute the indicator. We found, however, that our indicator improves significantly if the generated texts resemble real news articles in terms of length, style, structure etc. These findings are in line with Gao et al. (2022).
@@ A.2 Correlations with GDP growth
Tables A3A4 and A5 benchmark NEOS and its timelier versions against other survey-based indicators for economic outlook, and other text-based approaches introduced in Appendix A.3. The tables show the average correlations of the indicators with real year-on-year (yoy) and quarter-on-quarter (qoq) GDP growth for different lags over the periods Q1 1999 until Q1 2025, Q1 2009 until Q1 2025 and Q1 2014 until Q1 2025. Since GDP is a quarterly time series, we transform all monthly indicators to quarterly time series by taking three-month averages over the respective quarter. For instance, a lag of three quarters means that we compare GDP growth at time $t$ with the respective indicator at time t3t-3.
NEOS correlates with GDP growth in a similar way to the other indicators and outperforms for certain lags. The correlations with quarter-on-quarter GDP growth are much lower for all indicators. [^7] If qoq correlations with GDP growth are positive over multiple lags, yoy correlations will be higher by construction. Yoy GDP growth contains not only contemporaneous information but also information up to 12 months ago. NEOS shows the highest correlation for a lag of one quarter with respect to year-on-year GDP. The contemporaneous correlations (lag 0) can still be interpreted as a lead of the indicator since GDP is released with a delay of two months. [^8] The flash estimate is delayed by less than two months ($t+45$ days.) Note that Tables A3A4 and A5 show ex post comparisons and do not account for actual data availability of the other indicators in the past and GDP revisions. Not all indicators are released immediately at the end of the month and historically, some indicators are only available quarterly.
Table A3: Analysis from Q1 1999 until Q1 2025
@@
NEOS 0.60 0.68 0.64 0.51 0.24 0.64 0.53
NEOS (first 7 days in the month) 0.66 0.65 0.59 0.42 0.14 0.65 0.49
NEOS (first 14 days in the month) 0.66 0.66 0.60 0.43 0.14 0.66 0.50
NEOS (first 21 days in the month) 0.65 0.66 0.61 0.44 0.15 0.66 0.50
Lexicon-based approach 0.57 0.57 0.48 0.37 0.15 0.57 0.43
Manufacturing PMI for Switzerland 0.74 0.74 0.62 0.41 0.14 0.74 0.53
SECO Consumer Sentiment Index 0.46 0.41 0.35 0.22 -0.01 0.43 0.29
@@
NEOS 0.46 0.27 0.11 0.13 -0.07 0.36 0.18
NEOS (first 7 days in the month) 0.45 0.11 0.17 0.06 -0.08 0.28 0.14
NEOS (first 14 days in the month) 0.45 0.12 0.17 0.06 -0.08 0.29 0.14
NEOS (first 21 days in the month) 0.46 0.14 0.17 0.07 -0.08 0.30 0.15
Lexicon-based approach 0.37 0.19 0.01 0.10 -0.03 0.28 0.13
Manufacturing PMI for Switzerland 0.43 0.20 0.10 0.03 -0.08 0.32 0.14
SECO Consumer Sentiment Index 0.32 0.02 0.06 -0.00 -0.10 0.17 0.06
Correlations with GDP growth for different lags. The highest average correlations are written in bold.
@@
Table A4: Analysis from Q1 2009 until Q1 2025
lag of indicator (quarters) avg. over lags
Correlations with yoy GDP 0 1 2 3 4 $0,1$ $0,\ldots,4$
NEOS 0.48 0.50 0.42 0.30 -0.03 0.49 0.33
NEOS (first 7 days in the month) 0.56 0.47 0.38 0.18 -0.15 0.52 0.29
NEOS (first 14 days in the month) 0.55 0.48 0.40 0.20 -0.14 0.52 0.30
NEOS (first 21 days in the month) 0.55 0.49 0.40 0.22 -0.13 0.52 0.30
Lexicon-based approach 0.42 0.37 0.22 0.12 -0.12 0.40 0.20
KOF Business Situation Indicator 0.63 0.22 -0.05 -0.37 -0.63 0.42 -0.04
Manufacturing PMI for Switzerland 0.68 0.66 0.52 0.33 0.05 0.67 0.45
SECO Consumer Sentiment Index 0.37 0.29 0.23 0.11 -0.11 0.33 0.18
Correlations with qoq GDP
NEOS 0.36 0.08 -0.05 0.09 -0.19 0.22 0.06
NEOS (first 7 days in the month) 0.35 -0.14 0.11 -0.03 -0.19 0.11 0.02
NEOS (first 14 days in the month) 0.36 -0.12 0.10 -0.02 -0.19 0.12 0.02
NEOS (first 21 days in the month) 0.36 -0.10 0.09 -0.00 -0.20 0.13 0.03
Lexicon-based approach 0.23 0.04 -0.14 0.04 -0.13 0.14 0.01
KOF Business Situation Indicator 0.14 -0.38 -0.17 -0.21 -0.26 -0.12 -0.18
Manufacturing PMI for Switzerland 0.33 0.10 0.05 0.03 -0.12 0.22 0.08
SECO Consumer Sentiment Index 0.27 -0.12 0.02 -0.01 -0.08 0.07 0.02
Correlations with GDP growth for different lags. The highest average correlations are written in bold.
Table A5: Analysis from Q1 2014 until Q1 2025
lag of indicator (quarters) avg. over lags
Correlations with yoy GDP 0 1 2 3 4 $0,1$ $0,\ldots,4$
NEOS 0.48 0.52 0.47 0.38 -0.03 0.50 0.37
NEOS (first 7 days in the month) 0.56 0.46 0.44 0.26 -0.18 0.51 0.31
NEOS (first 14 days in the month) 0.55 0.47 0.45 0.27 -0.17 0.51 0.32
NEOS (first 21 days in the month) 0.55 0.48 0.46 0.29 -0.15 0.51 0.33
Lexicon-based approach 0.43 0.40 0.26 0.17 -0.13 0.41 0.23
KOF Business Situation Indicator 0.63 0.19 -0.03 -0.35 -0.65 0.41 -0.04
Manufacturing PMI for Switzerland 0.67 0.64 0.54 0.40 0.11 0.66 0.47
Service PMI for Switzerland 0.67 0.48 0.29 0.16 -0.32 0.58 0.26
SECO Consumer Sentiment Index 0.39 0.26 0.22 0.13 -0.11 0.33 0.18
Correlations with qoq GDP
NEOS 0.43 0.09 -0.06 0.14 -0.23 0.26 0.07
NEOS (first 7 days in the month) 0.44 -0.19 0.16 -0.01 -0.25 0.12 0.03
NEOS (first 14 days in the month) 0.44 -0.17 0.14 -0.06 -0.26 0.14 0.03
NEOS (first 21 days in the month) 0.44 -0.14 0.13 -0.07 -0.26 0.15 0.04
Lexicon-based approach 0.29 0.06 -0.18 0.07 -0.15 0.18 0.02
KOF Business Situation Indicator 0.18 -0.41 -0.15 -0.20 -0.27 -0.12 -0.17
Manufacturing PMI for Switzerland 0.36 0.12 0.07 0.07 -0.11 0.24 0.10
Service PMI for Switzerland 0.50 -0.09 -0.15 -0.04 -0.25 0.20 -0.01
SECO Consumer Sentiment Index 0.29 -0.14 0.04 0.00 -0.07 0.08 0.02
Correlations with GDP growth for different lags. The highest average correlations are written in bold.
@@
A.3 EPU and lexicon-based approach to measure sentiment from news articles
We compute two alternative text-based indicators, which are based on the same underlying news articles data as NEOS.
a) We closely follow Baker et al. (2016) to construct an EPU index for Switzerland, using their keywords plus Switzerland all translated to German and French.